CASH CONVERSION CYCLE

Chastain Corporation is trying to determine the effect of its inventory turnover ratio and days sales outstanding (DSO) on its cash conversion cycle. Chastain's 2016 sales (all on credit) were $272,000; its cost of goods sold is 80% of sales; and it earned a net profit of 3%, or $8,160. It turned over its inventory 4 times during the year, and its DSO was 35 days. The firm had fixed assets totaling $32,000. Chastain's payables deferral period is 35 days. Assume 365 days in year for your calculations.

  1. Calculate Chastain's cash conversion cycle. Round your answer to two decimal places. Do not round intermediate calculations.
      days

  2. Assuming Chastain holds negligible amounts of cash and marketable securities, calculate its total assets turnover and ROA. Round your answers to two decimal places. Do not round intermediate calculations.
    Total assets turnover   
    ROA %

  3. Suppose Chastain's managers believe that the inventory turnover can be raised to 9.9 times. What would Chastain's cash conversion cycle, total assets turnover, and ROA have been if the inventory turnover had been 9.9 for 2016? Round your answers to two decimal places. Do not round intermediate calculations.
    Cash conversion cycle days
    Total assets turnover   
    ROA %
Answer #1

Answer a.

Days Sales Outstanding = 35 days

Days Inventory Outstanding = 365 / Inventory Turnover
Days Inventory Outstanding = 365 / 4
Days Inventory Outstanding = 91.25 days

Payable Deferral Period = 35 days

Cash Conversion Cycle = Days Sales Outstanding + Days Inventory Outstanding - Payable Deferral Period
Cash Conversion Cycle = 35 + 91.25 - 35
Cash Conversion Cycle = 91.25 days

Answer b.

Days Sales Outstanding = 365 * Accounts Receivable / Sales
35 = 365 * Accounts Receivable / $272,000
Accounts Receivable = $26,082.19

Cost of Goods Sold = 80% * Annual Sales
Cost of Goods Sold = 80% * $272,000
Cost of Goods Sold = $217,600

Inventory Turnover = Cost of Goods Sold / Inventory
4 = $217,600 / Inventory
Inventory = $54,400

Total Assets = Accounts Receivable + Inventory + Fixed Assets
Total Assets = $26,082.19 + $54,400 + $32,000
Total Assets = $112,482.19

Net Income = Sales * Profit Margin
Net Income = $272,000 * 3%
Net Income = $8,160

Total Assets Turnover = Sales / Total Assets
Total Assets Turnover = $272,000 / $112,482.19
Total Assets Turnover = 2.42 times

Return on Assets = Net Income / Total Assets
Return on Assets = $8,160 / $112,482.19
Return on Assets = 7.25%

Answer c.

Days Sales Outstanding = 35 days

Days Inventory Outstanding = 365 / Inventory Turnover
Days Inventory Outstanding = 365 / 9.90
Days Inventory Outstanding = 36.87 days

Payable Deferral Period = 35 days

Cash Conversion Cycle = Days Sales Outstanding + Days Inventory Outstanding - Payable Deferral Period
Cash Conversion Cycle = 35 + 36.87 - 35
Cash Conversion Cycle = 36.87 days

Days Sales Outstanding = 365 * Accounts Receivable / Sales
35 = 365 * Accounts Receivable / $272,000
Accounts Receivable = $26,082.19

Cost of Goods Sold = 80% * Annual Sales
Cost of Goods Sold = 80% * $272,000
Cost of Goods Sold = $217,600

Inventory Turnover = Cost of Goods Sold / Inventory
9.90 = $217,600 / Inventory
Inventory = $21,979.80

Total Assets = Accounts Receivable + Inventory + Fixed Assets
Total Assets = $26,082.19 + $21,979.80 + $32,000
Total Assets = $80,061.99

Net Income = Sales * Profit Margin
Net Income = $272,000 * 3%
Net Income = $8,160

Total Assets Turnover = Sales / Total Assets
Total Assets Turnover = $272,000 / $80,061.99
Total Assets Turnover = 3.40 times

Return on Assets = Net Income / Total Assets
Return on Assets = $8,160 / $80,061.99
Return on Assets = 10.19%